# Statistical signatures of quantum contextuality

### v1

@@ I Introduction
The interpretation of quantum mechanics is controversial because it is impossible to identify individual measurement outcomes with an underlying measurement independent reality [1, 2]. Unfortunately, the familiar terminology of quantum mechanics tends to encourage images of underlying realities by carelessly identifying quantum state components with measurement outcomes even when no such measurements are ever performed. It cannot be stressed enough that quantum mechanics contradicts such fantasies directly by explaining the relation between different measurement contexts in terms of superpositions of their outcomes. Still, quantum mechanics does not seem to provide any good alternative. How can it be that we can describe the failure of objective realism in great detail, but cannot find any good alternative that would explain the relation beteen different measurement contexts in a manner consistent with the highly successful theoretical formalism? Perhaps we need a more appropriate formulation of the problem, so that we can identify the correct relations between different measurements without any speculation about unseen realities.
@@
Conventional formulations of quantum mechanics tend to prefer a specific context in the formulation of quantum states. To better understand the origin of contextuality in the quantum formalism, we need to investigate statistical expressions that relate different measurement contexts to each other. One such approach is to interpret the quantum formalism as a generalized probability theory [10, 11, 12, 13]. Not surprisingly, the results show that the quantum formalism introduces non-positive quasi probabilities into the statistics, which explains the differences between the predictions of quantum mechanics and the non-contextual hidden variable theories proposed as an alternative. The problem is that there is a wide range of possible quasi probabilities based on different selections of measurements or different operator expansions [14]. It may therefore be more promising to investigate the relation between quantum states and experimentally observable statistics in more detail. Here, I will take a look at the possible reconstruction of quantum states from experimentally accessible data, also known as quantum state tomography. Originally, these methods emphasized the similarity between quantum theory and classical statistics, e.g. when a semi classical analysis is used to reconstruct a phase space representation of the quantum state [15]. It has also been shown that weak measurements or a variety of joint measurements can be used to reconstruct a Kirkwood-Dirac distribution of two non-commuting observables [16, 17, 18, 19, 20, 21]. These approaches to quantum tomography show that the reconstruction of an unknown state can be achieved by applying methods of classical statistics to characterize the non-classical correlations between the outcomes of incompatible measurements. Non-positive quasi probabilities are obtained because these correlations cannot be explained by a positive valued probability distribution over the possible combination of incompatible outcomes. As observed very early in the history of quantum mechanics, any attempt to describe quantum states as joint probabilities of complementary observables necessarily results in such non-positive distributions [22, 23].
@@
Quantum theory makes precise statements about the relations between different contexts. These statements show that quantum mechanics is not constraint by the assumptions of a measurement independent reality. Instead, the uncertainties of the initial state obtain a different meaning in each of the different contexts. This is not just an interpretation of the formalism, the universality of tomographic relations indicate that these are fundamental laws of physics. By designing experiments with sufficient care, it is possible to characterize the specific dependence of reality on the type of meaurement that will eventually be performed. The relations presented in the following are a first step towards a fully deterministic description of the physical world that resolves the measurement problem by identifying the precise role that the measurement context plays in shaping the observable reality of all objects. It should be noted that this kind of progress is made possible by the realization that there is no need for a measurement independent reality. The application of the formalism strongly suggests that the familiar structure of reality that we experience in our actual lives is sufficiently explained in terms of the deterministic relations between different measurements. This is a point that is often overlooked in discussions of the measurement problem such as [1, 2]. We do not have any direct access to any microscopic realities, and our ability to perform measurements is limited by our macroscopic interventions in the world around us. If we cannot reconcile our actual experience with our descriptions of “reality,” we need to question the description and not the experience.
@@ II Characterization of measurement contexts
Many fundamental investigations of quantum statistics start with the two dimensional Hilbert space of a logical qubit. The physical analogy is a spin-1/2 system, where every direction in space is a potential measurement basis. It is difficult to formulate a logical relation between the different measurement contexts, since no two measurements can be performed jointly and each measurement outcome belongs to only one possible measurement context. The situation changes when a three dimensional Hilbert space is considered. It is now possible to construct a much wider range of possible measurements, and two different measurement contexts can share the same measurement outcome. The Kochen-Specker theorem uses such relations to investigate the statistical relations between incompatible measurements, demonstrating that the relation between different measurement contexts cannot be described by uncertainty limited classical statistics [5]. In the most simple scenario, five different measurement contexts are related to each other by shared measurement outcomes in a three dimensional Hilbert space [6, 7, 8, 9]. As shown in figure 1, the symmetry of this scenario can be illustrated by a pentagram, where each measurement context is represented by a triangle and the shared outcomes form a pentagon at the center. Note that this representation is a more complete version of the graph derived in [24]. I am using the same terminology, with the context $\{1,2,3\}$ at the top, the contexts $\{1,S1,D1\}$ and $\{2,S2,D2\}$ sharing the outcomes $1$ and $2$ at their respective sides, and the contexts $\{S1,f,P1\}$ and $\{S2,f,P2\}$ at the bottom, joined by the shared outcome $f$.
@@
What can be said about the relation between different measurement contexts? The identification of measurement contexts with the different stages of single-particle propagation in a three-path interferometer illustrates this relation by associating measurement outcomes with physical paths. A non-contextual hidden variable model would then assign a specific path through the interferometer to each particle. Since the input context $\{1,2,3\}$ is perfectly recovered at the output ports, paths originating from input $1$ must end up in output $1$ and so forth, leaving a total of eleven distinct possibilities for the propagation of the particle. Table 1 illustrates the eleven paths by assigning a sequence of Hilbert space vectors to each of them. As the table shows, two path segments $[a,b]$ are sufficient to define each of the paths, except for the path composed of the five segments that are unique to each context, which is therefore labeled $[0]$. The quantum mechanical relations that replace the sequence of paths in the Hilbert space formalism are inner products of the state vectors that represent the measurement outcomes of the different contexts [24]. Although quantum contextuality shows that quantum states cannot be represented by positive valued probability distributions over the eleven paths in table 1, it should be possible to find a quasi-probability representation of the density matrix that relates directly to these $11$ paths. It has already been noted elsewhere that Kirkwood-Dirac quasiprobabilities are ideally suited for this purpose [25, 26, 27]. In the following, I will develop an expansion of the density matrix that can serve as a starting point for a detailed investigation of the relations between the five measurement contexts described by the Hilbert space formalism.
@@ III Relations between non-orthogonal states
As explained in previous work, Kirkwood-Dirac distributions are the most consistent representation of quantum statistics, accurately reproducing the relations between non-orthogonal states in the Hilbert space formalism [28, 29, 30, 31, 32, 33, 34, 35]. Such relations represent the non-classical structure of operator statistics and ensure that quantum theory cannot be reconciled with any classical statistical theory [14, 31]. Here, the goal is to identify a specific set of relations between five different measurement contexts that make optimal use of the orthogonality relations used to demonstarte quantum contextuality. To achieve this, we can start with a selection of non-orthogonal states representing one of the five trajectories composed of three states, e.g. $\{\left|1\right\rangle,\left|f\right\rangle,\left|D2\right\rangle\}$. These three state vectors form a complete non-orthogonal basis of the three-dimensional Hilbert space. The corresponding set of contravariant vectors is $\{\left|S2\right\rangle,\left|2\right\rangle,\left|S1\right\rangle\}$, and the expansion of any state vector in this non-orthogonal basis can be expressed by representing the identity operator as a sum of three non-hermitian operators,
@@ IV Complete expansions of arbitrary states
As mentioned before, the Kirkwood-Dirac quasi probabilities describe the relation between the quantum statistics of $\left|a\right\rangle$ and $\left|b\right\rangle$ in terms of the weak values of $\left|a\right\rangle\hskip-2.84526pt\left\langle a\right|$ post-selected in $\left|b\right\rangle$ [16, 17]. In the following, I will refer to them as KD terms. When the non-orthogonal expansion of Eq.(6) is applied to a density operator $\hat{\rho}$, the coefficients of the expansion can be expressed by a combination of probabilities and KD terms,

### v2

@@ I Introduction
The interpretation of quantum mechanics is controversial because it is impossible to identify individual measurement outcomes with an underlying measurement independent reality [1, 2]. Unfortunately, the familiar terminology of quantum mechanics tends to encourage images of underlying realities by carelessly identifying quantum state components with measurement outcomes even when no such measurements are ever performed. It cannot be stressed enough that quantum mechanics contradicts such fantasies directly by explaining the relation between different measurement contexts in terms of superpositions of their outcomes. Still, quantum mechanics does not seem to provide any good alternative. How can it be that we can describe the failure of objective realism in great detail, but cannot find any good alternative that would explain the relation between different measurement contexts in a manner consistent with the highly successful theoretical formalism? Perhaps we need a more appropriate formulation of the problem, so that we can identify the correct relations between different measurements without any speculation about unseen realities.
@@
Conventional formulations of quantum mechanics tend to represent states in a specific context, effectively hiding the relations with other contexts in the coherences between orthogonal states that represent the different outcomes of a single measurement context. In fact, the origin of contextuality can always be traced to state independent relations between different contexts expressed in terms of fundamental relations between non-commuting operators [10, 11, 12, 13]. Nevertheless, the experimentally observable evidence can only be obtained in the form of measurement statistics associated with specific quantum states [14]. In the end, contextuality describes statistical relations between different measurements that apply to arbitrary quantum states. To better understand the origin of contextuality in the quantum formalism, it is therefore useful to investigate representations of quantum states that relate different measurement contexts to each other. This can be done by interpreting the quantum formalism as a generalized probability theory [15, 16, 17, 18]. Not surprisingly, the results show that the quantum formalism introduces non-positive quasi probabilities into the statistics, which explains the differences between the predictions of quantum mechanics and the non-contextual hidden variable theories proposed as an alternative. The problem is that there is a wide range of possible quasi probabilities based on different selections of measurements or different operator expansions [19]. To identify an appropriate quasi probability, it is necessary to consider the specific relations between the different measurement contexts that characterize any given scenario. Here, I will take a look at the possible reconstruction of quantum states from experimentally accessible data, also known as quantum state tomography. Originally, these methods emphasized the similarity between quantum theory and classical statistics, e.g. when a semi classical analysis is used to reconstruct a phase space representation of the quantum state [20]. It has also been shown that weak measurements or a variety of joint measurements can be used to reconstruct a Kirkwood-Dirac distribution of two non-commuting observables [21, 22, 23, 24, 25, 26]. These approaches to quantum tomography show that the reconstruction of an unknown state can be achieved by applying methods of classical statistics to characterize the non-classical correlations between the outcomes of incompatible measurements. Non-positive quasi probabilities are obtained because these correlations cannot be explained by a positive valued probability distribution over the possible combination of incompatible outcomes. As observed very early in the history of quantum mechanics, any attempt to describe quantum states as joint probabilities of complementary observables necessarily results in such non-positive distributions [27, 28].
@@
Quantum theory makes precise statements about the relations between different contexts. These statements show that quantum mechanics is not constrained by the assumptions of a measurement independent reality. Instead, the uncertainties of the initial state obtain a different meaning in each of the different contexts. This is not just an interpretation of the formalism, the universality of tomographic relations indicate that these are fundamental laws of physics. By designing experiments with sufficient care, it is possible to characterize the specific dependence of reality on the type of measurement that will eventually be performed. The relations presented in the following are a first step towards a fully deterministic description of the physical world that resolves the measurement problem by identifying the precise role that the measurement context plays in shaping the observable reality of all objects. It should be noted that this kind of progress is made possible by the realization that there is no need for a measurement independent reality. The application of the formalism strongly suggests that the familiar structure of reality that we experience in our actual lives is sufficiently explained in terms of the deterministic relations between different measurements. This is a point that is often overlooked in discussions of the measurement problem such as [1, 2]. We do not have any direct access to any microscopic realities, and our ability to perform measurements is limited by our macroscopic interventions in the world around us. If we cannot reconcile our actual experience with our descriptions of “reality,” we need to question the description and not the experience.
@@ II Characterization of measurement contexts
Many fundamental investigations of quantum statistics start with the two dimensional Hilbert space of a logical qubit. The physical analogy is a spin-1/2 system, where every direction in space is a potential measurement basis. It is difficult to formulate a logical relation between the different measurement contexts, since no two measurements can be performed jointly and each measurement outcome belongs to only one possible measurement context. The situation changes when a three dimensional Hilbert space is considered. It is now possible to construct a much wider range of possible measurements, and two different measurement contexts can share the same measurement outcome. The Kochen-Specker theorem uses such relations to investigate the statistical relations between incompatible measurements, demonstrating that the relation between different measurement contexts cannot be described by uncertainty limited classical statistics [5]. In the most simple scenario, five different measurement contexts are related to each other by shared measurement outcomes in a three dimensional Hilbert space [6, 7, 8, 9]. Each measurement context is described by an orthogonal basis, where one of the basis states is unique to this context and the other two are shared with two other contexts, resulting in a cyclic relation between all five contexts. As shown in figure 1, the symmetry of this scenario can be illustrated by a pentagram, where each measurement context is represented by a triangle and the shared outcomes form a pentagon at the center. Note that this representation is a more complete version of the graph derived in [29]. I am using the same terminology, with the context $\{1,2,3\}$ at the top, the contexts $\{1,S1,D1\}$ and $\{2,S2,D2\}$ sharing the outcomes $1$ and $2$ at their respective sides, and the contexts $\{S1,f,P1\}$ and $\{S2,f,P2\}$ at the bottom, joined by the shared outcome $f$. A statstical proof of contextuality can now be formulated using only the five shared outcomes. Since each of these five outcomes excludes two of the others, non-contextual logic requires that no more than two of the shared outcomes can be true at the same time. The sum of the probabilities for these five outcomes should therefore have an upper bound of two. However, this restriction does not apply in Hilbert space, resulting in possible violations of the non-contextual bound.
@@
What can be said about the relation between different measurement contexts? The identification of measurement contexts with the different stages of single-particle propagation in a three-path interferometer illustrates this relation by associating measurement outcomes with physical paths. A non-contextual hidden variable model would then assign a specific path through the interferometer to each particle. Since the input context $\{1,2,3\}$ is perfectly recovered at the output ports, paths originating from input $1$ must end up in output $1$ and so forth, leaving a total of eleven distinct possibilities for the propagation of the particle. Table 1 illustrates the eleven paths by assigning a sequence of Hilbert space vectors to each of them. As the table shows, two path segments $[a,b]$ are sufficient to define each of the paths, except for the path composed of the five segments that are unique to each context, which is therefore labeled $[0]$. The quantum mechanical relations that replace the sequence of paths in the Hilbert space formalism are inner products of the state vectors that represent the measurement outcomes of the different contexts [29]. Although quantum contextuality shows that quantum states cannot be represented by positive valued probability distributions over the eleven paths in table 1, it should be possible to find a quasi probability representation of the density matrix that relates directly to these $11$ paths. It has already been noted elsewhere that Kirkwood-Dirac quasi probabilities are ideally suited for this purpose [30, 31, 32]. In the following, I will develop an expansion of the density matrix that can serve as a starting point for a detailed investigation of the relations between the five measurement contexts described by the Hilbert space formalism.
@@ III Relations between non-orthogonal states
As explained in previous work, Kirkwood-Dirac distributions are the most consistent representation of quantum statistics, accurately reproducing the relations between non-orthogonal states in the Hilbert space formalism [33, 34, 35, 36, 37, 38, 39, 40, 41]. Such relations represent the non-classical structure of operator statistics and ensure that quantum theory cannot be reconciled with any classical statistical theory [19, 36]. Here, the goal is to identify a specific set of relations between five different measurement contexts that make optimal use of the orthogonality relations used to demonstarte quantum contextuality. To achieve this, we can start with a selection of non-orthogonal states representing one of the five trajectories composed of three states, e.g. $\{\left|1\right\rangle,\left|f\right\rangle,\left|D2\right\rangle\}$. These three state vectors form a complete non-orthogonal basis of the three-dimensional Hilbert space. The corresponding set of contravariant vectors is $\{\left|S2\right\rangle,\left|2\right\rangle,\left|S1\right\rangle\}$, and the expansion of any state vector in this non-orthogonal basis can be expressed by representing the identity operator as a sum of three non-hermitian operators,
@@ IV Complete expansions of arbitrary states
As mentioned before, the Kirkwood-Dirac quasi probabilities describe the relation between the quantum statistics of $\left|a\right\rangle$ and $\left|b\right\rangle$ in terms of the weak values of $\left|a\right\rangle\hskip-2.84526pt\left\langle a\right|$ post-selected in $\left|b\right\rangle$ [21, 22, 43]. In the following, I will refer to them as KD terms. When the non-orthogonal expansion of Eq.(6) is applied to a density operator $\hat{\rho}$, the coefficients of the expansion can be expressed by a combination of probabilities and KD terms,